Nitric acid is often manufactured from the atmospheric gases nitrogen and oxygen, and hydrogen prepared by reforming natural gas, in a two-step process. In the first step, nitrogen and hydrogen react to form ammonia: N2+3H2=2NH3

In the second step, ammonia and oxygen react to form nitric acid HNO3 and water: NH3+ 2O2=HNO3+H2O

Suppose the yield of the first step is 96.% and the yield of the second step is 85.%. Calculate the mass of hydrogen required to make 5.0kg of nitric acid.

Be sure your answer has a unit symbol, if needed, and is rounded to 2 significant digits.
Answer #1

1) N2 + 3H2 = 2NH3

2) NH3 + 2O2 = HNO3 + H2O

no of mol of HNO3 = 5/63 = 0.08 kmol

no of mol of NH3 required = 0.08*100/85 = 0.094 kmol

eq2 , 3 mol H2 = 2 mol NH3

no of mol of H2 required = 0.094*(3/2)*(100/96) =

                         = 0.147 kmol

mass of H2 required = 0.147*2 = 0.294 kg

                    = 0.29 kg

answer: mass of H2 required = 0.29 kg